# Underwater Image Enhancement by Diffusion Model with Customized CLIP-Classifier

### v1

@@ Abstract
In this paper, we propose a novel underwater image enhancement method, by utilizing the multi-guided diffusion model for iterative enhancement. Unlike other image enhancement tasks, underwater images suffer from the unavailability of real reference images. Although existing works exploit synthetic images, manually selected well-enhanced images as reference images, to train enhancement networks, their enhancement performance always comes with subjective preferences that are inherited from the manual selection. To address this issue, we also use the image synthesis strategy, but the synthetic images derive from in-air natural images degraded into corresponding underwater images, guided by the underwater domain. Based on this strategy, the diffusion model can learn the prior knowledge of image enhancement from the underwater degradation domain to the real in-air natural domain. However, it is inevitable to fine-tune the model to suit downstream tasks, and this may erase the prior knowledge. To mitigate this, we combine the prior knowledge from the in-air natural domain with Contrastive Language-Image Pretraining (CLIP) to train a classifier for controlling the diffusion model generation process. Moreover, for image enhancement tasks, we find that the image-to-image diffusion model and the CLIP-Classifier mainly act in the high-frequency region during the fine-tuning process. Therefore, we propose a fast fine-tuning strategy focusing on the high-frequency region, which can be up to 10 times faster than the traditional strategy. Extensive experiments demonstrate that our method, abbreviated as CLIP-UIE, exhibit a more natural appearance. The source code and pre-trained models are available on the project homepage: https://github.com/OUCVisionGroup/CLIP-UIE.
@@ I Introduction
Underwater image enhancement (UIE) is the process of generating a visually clear image that is consistent with an input-degraded underwater image, which often exhibits color casts, low contrast, and poor visibility [1, 2, 3]. It plays a significant role in vision-based research for marine and underwater tasks, e.g. marine animal detection [4], ecosystem protection [5], and underwater robot application [6], etc.
@@
By reviewing typical UIE strategies, early approaches, based on the physical model, focus on the accurate estimation of medium transmission that represents the percentage of the scene radiance reaching the camera [7, 8, 9]. Though physical model-based methods can achieve promising performance in some cases, they tend to produce unnatural and poor results in challenging underwater scenarios [10]. Recent approaches based on deep learning technology have gradually outperformed traditional physical model-based approaches. [11, 12, 13]. While simple regression-based methods with feedback convolutional networks may work for image enhancement in simple underwater scenarios, they often lacks generalisation and robustness for complex scenes as they cannot fully model the real-underwater degradation. The image-to-image diffusion models have been successful in learning complex empirical distributions of images, enabling the transformation of images between different domains through iterative refinement steps.
@@
However, image-to-image diffusion models heavily rely on paired datasets, and the inability to capture clear real-world reference images poses a challenge for UIE [14, 15, 16]. To address this issue, researchers employed various approaches to synthesize multiple enhanced images of a single image, and then manually select the visually optimal one as reference [17, 18]. Nevertheless, the synthesized paired datasets contain subjective preferences, reflected in the enhanced results, resulting in natural domain shift and weakened generalization. As for the unsupervised or semi-surprised strategies, they can also generate pleasing underwater images [19, 18], without paired datasets or relying on limited paired samples, but the performance is not superior to the fully supervised method. The other alternative approach is to combine a pre-trained diffusion model, which captures the prior knowledge of the mapping transitions between different image domains, with custom datasets, thus reducing dependency on paired datasets [20]. However, the fine-tuning, to directly continue training with the additional paired data, may destroy the valuable prior knowledge [21]. Moreover, another limitation, the quality of the enhancement result is limited by the reference domain due to the lack of real image guidance, resulting in unnatural and unrealistic enhanced images.
@@
To address the above limitations, we propose CLIP-UIE, a new approach to treat UIE as a conditional underwater image generation, inspired by recent works on Denoising Diffusion Probabilistic Models [22, 23, 24, 25] and CLIP guidance for image manipulation [26, 27, 28]. The proposed method is trained with a denoising objective to iterative refinement enhancement images by giving a degraded image. Initially, the key is to obtain a large number of underwater paired datasets. We utilize the image synthesis strategy based on color transfer [29, 30] to degrade in-air natural images into the real underwater domain, synthesising the corresponding degraded underwater images, which is free of any subjective preferences. Subsequently, we train the basic image-to-image diffusion model on this synthetic paired dataset to learn the prior knowledge of the mapping transitions from the underwater degradation domain to real in-air natural domain. This prior knowledge can adverse impact of unrealistic and unnatural synthetics effectively. However, it is inevitable to fine-tune the model to suit downstream tasks, and this may erase the prior knowledge, leading to overfitting and mode collapse [21]. Song et al. [31] and Jonathan et al. [32] find that a pre-trained diffusion model can be conditioned using the gradients of a classifier. Thus, we can exploit a trained classifier with prior knowledge, combined with custom datasets to jointly fine-tune the diffusion model to avoid overfitting and mode collapse, bridging the domain gap between synthetic and real images. It has been shown that Vision-Language models encapsulate generic information, that can be paired with generative model to provide a simple and intuitive text-driven interface for image generation and manipulation [33], particularly CLIP. Although CLIP can serve as a classifier to distinguish in-air natural and underwater degraded images to a certain extent, using its gradients directly to control the reverse diffusion process is still non-trivial. Moreover, it is barely possible to find accurate prompts to describe the types of underwater degradation, and rewriting similar prompts easily leads to huge differences in CLIP score, as shown in Fig. 1. We train a CLIP-Classifier that contains prior knowledge of the in-air natural domain by constraining text-image similarity in the CLIP embedding space. Moreover, for image enhancement tasks, we find that the image-to-image diffusion model and the CLIP-Classifier mainly act in the high-frequency region during the fine-tuning process. Therefore, we propose a fast fine-tuning strategy that focuses on the high-frequency range and can be up to 10 times faster. The contributions of this paper are summarized as follows:
@@
  • We propose to use color transfer to degrade 500k in-air natural images from INaturalist [34] into underwater images guided by the real underwater domain, overcoming subjective preferences introduced by manual selection of reference images. The pre-trained image-to-image diffusion model is trained from scratch on this synthetic dataset.
@@
  • We propose a CLIP-Classifier that inherits prior knowledge of the in-air natural domain, and then combine it with custom datasets to jointly fine-tune the diffusion model to mitigate catastrophic forgetting and mode collapse. Experiments and ablation studies verify the good performance of the proposed CLIP-UIE and it breaks the limitations of the reference domain to a certain extent.
@@
  • We find that for image enhancement tasks, which require consistent content, the image-to-image diffusion model and the CLIP-Classifier mainly act in high-frequency regions. Therefore, we propose a new fine-tuning strategy that acts only on high-frequency regions, which significantly improves the fine-tuning speed, even up to 10 times.
@@
Figure 1: CLIP with proper prompts can serve as a classifier to evaluate the image quality and distinguish between in-air natural and underwater images. However, our simple approximate rewriting of the prompt results in a dramatic change in CLIP score. On the contrary, our CLIP-Classifier with learned prompt shows more robust results for different types of in-air natural/underwater scenes.
@@ II-A Underwater Image Enhancement
Several approaches have been proposed in the underwater enhancement image, categorizing into two groups: traditional methods and deep learning-based methods. For the traditional methods, researchers focus on directly adjusting image pixel values to produce a subjectively and visually appealing image, using techniques such as image fusion [35], histogram equalization [36], and pixel distribution [37], etc. Ancuti et al. [38] propose the fusion-based underwater image enhancement method, exploiting two images derived from a color-compensated and white-balanced version to prompt color contrast. However, these traditional methods are not sensitive to the types of underwater degradation, due to the omission of the underwater imaging mechanism. To improve the visual quality, physical model-based methods directly estimate the parameters of underwater imaging models based on prior information. For example, Fu et al. [7] utilize the retinex-based approach to capture a clean underwater image by combining enhanced reflectance and illumination. To further promote the performance, Zhuang et al. [9] improve the retinex-based algorithm by introducing Bayesian to estimate underwater medium transmissions. Nevertheless, owing to the scarcity of prior information, the enhanced image easily becomes unnatural artifacts and over/under-enhanced.
@@
Recently, deep learning has made significant advances in computer vision and image processing, there are serial attempts made to improve the performance of underwater image enhancement by data-driven strategy. However, underwater images suffer from the unavailability of real reference images, leading researchers to synthesize paired datasets. Concretely, Li et al. [17] construct an underwater image benchmark (UIEB) by selecting the best result of various approaches as the reference. Nevertheless, the synthetic paired datasets cannot avoid subjective preferences from the manual selection. To relax the requirement of paired underwater images, several GAN-based weakly supervised methods can generate pleasing underwater images [19, 18], but the performance is not superior to the fully supervised method.
@@
The strategy of using in-air natural images to synthesise paired underwater degradation datasets provides a good solution. Du et al. [20] transform in-air natural images into the pseudo underwater image domain by Koschmieder’s light scattering model, utilizing the prior knowledge of in-air natural images to counteract the adverse impact of unreal synthetic. However, the pseudo-underwater domain based on the light scattering model cannot fully model the complexity of real-world underwater degradation. To solve this issue, we propose a new synthesis strategy based on color transfer [29] to yield paired datasets, degrading in-air natural images to real underwater domain. By this means, we introduce more prior knowledge: the mapping between the real underwater degradation domain and the real in-air natural domain.
@@ II-B Diffusion Models
Diffusion models recently have impressive performance on image generation by inverting image noising process with reverse-time stochastic differential equations (SDEs) [39]. However, this inversion process tends to change the image context. To condition the denoising process, Classifier Guidance [32, 31] is an effective method to improve the sample quality of conditional diffusion models with the gradients of the classifier. An alternative strategy is to establish the relationship between input and conditional information in the model’s latent space [40, 25]. Controlling diffusion models facilitates personalisation, customization, or task-specific image generation. Hence, the diffusion models have been successfully applied to various image enhancement tasks. The image-to-image diffusion model was first introduced by Saharia et al. [25] and has been exploited for image enhancement [41, 20], inpainting [42], and super-resolution [25], etc.
Conditional diffusion models can learn the mapping between different image domains [21]. Based on this observation, our proposal uses the aforementioned color transfer to synthetic images, and trains the conditional diffusion model to learn the mapping between the real underwater degradation domain and the real in-air natural domain. However, this model will be guided by the latter condition without real-underwater reference, and this may destroy prior knowledge. To mitigate this issue, we train a classifier to discriminate between in-air natural and underwater degraded images, integrating the prior knowledge of the in-air natural domain into this classifier, and then combine it with custom datasets to jointly fine-tune the diffusion model, enabling the transfer of enhancement results to both the natural and reference domains.
@@ II-C CLIP-based Image Manipulation
The CLIP model benefits from 400 million curated image-text pairs can predict the correct pairings of a batch of (image, text) unseen examples, and consists of a text encoder and an image encoder by jointly training [43]. It has shown remarkable performance in text-to-image generation. Gal et al. [33] and Patashnik et al. [44] use CLIP models to modify StyleGAN-generated images with text prompts. In addition, Kim et al. [26] and Avrahami et al. [40] propose ways to utilize the diffusion model guided by CLIP models for global text-to-image synthesis, as well as local image editing. Text-driven generative models work well for objective things with clear “word” prompts, such as ‘apple’, ‘banana’, and ‘orange’, whereas it is barely possible to find accurate prompts for abstract image styles and image visual qualities. Moreover, prompt engineering for abstract things requires domain experts to annotate each image, and the more challenging is the similar prompts may get different CLIP score [28] (see Fig. 1). This limits the user to custom domain generation and manipulation for the abstract semantic concepts with the pre-trained CLIP.
To overcome the problems above, fine-tuning the CLIP model with considerable data for domain adaptation is a recent emerging research [33], but it is labor-intensive and time-consuming. Concurrently, prompt learning with frozen CLIP model is the alternation to extract accurate low-level image representations [28], such as luminance, exposure, and contrast [27]. We also use prompt learning to train a CLIP-Classifier, distinguishing between in-air natural and underwater degraded images. Then, the gradients of this CLIP-Classifier are used to guide the sampling process of the diffusion model.
Figure 2: The preparation for the pre-trained model. (a) Randomly select template A from the template pool (underwater domain). Then, the Color Transfer module, guided by template A, degrades a in-air natural image from INaturalist [34] into underwater domain, constructing paired datasets for training image-to-image diffusion model. (b) The image-to-image diffusion model SR3 [25] is trained to learn the prior knowledge, the mapping from the real underwater degradation domain and the real in-air natural domain, and to generate the corresponding enhancement results under the condition of the input synthetic underwater images produced by Color Transfer.
III Proposed Method
III-A Preliminary of Diffusion Models
Diffusion probabilistic models are a type of latent variable models that define a diffusion Markov chain of diffusion process to slowly inject random noise to the data, and a corresponding reverse process to slowly remove the noise [22]. The diffusion process is a forward process that gradually injects noise into the data sample ${x}_{0}$ over $T$ time steps, diffusing the data samples through Gaussian transition:

### v2

@@ Abstract
Underwater Image Enhancement (UIE) aims to improve the visual quality from a low-quality input. Unlike other image enhancement tasks, underwater images suffer from the unavailability of real reference images. Although existing works exploit synthetic images and manually select well-enhanced images as reference images to train enhancement networks, their upper performance bound is limited by the reference domain. To address this challenge, we propose CLIP-UIE, a novel framework that leverages the potential of Contrastive Language-Image Pretraining (CLIP) for the UIE task. Specifically, we propose employing color transfer to yield synthetic images by degrading in-air natural images into corresponding underwater images, guided by the real underwater domain. This approach enables the diffusion model to capture the prior knowledge of mapping transitions from the underwater degradation domain to the real in-air natural domain. Still, fine-tuning the diffusion model for specific downstream tasks is inevitable and may result in the loss of this prior knowledge. To migrate this drawback, we combine the prior knowledge of the in-air natural domain with CLIP to train a CLIP-Classifier. Subsequently, we integrate this CLIP-Classifier with UIE benchmark datasets to jointly fine-tune the diffusion model, guiding the enhancement results towards the in-air natural domain. Additionally, for image enhancement tasks, we observe that both the image-to-image diffusion model and CLIP-Classifier primarily focus on the high-frequency region during fine-tuning. Therefore, we propose a new fine-tuning strategy that specifically targets the high-frequency region, which can be up to 10 times faster than traditional strategies. Extensive experiments demonstrate that our method exhibits a more natural appearance. The source code and pre-trained models are available on the project homepage: https://oucvisiongroup.github.io/CLIP-UIE.html/.
@@ I Introduction
With the rise of marine ecological research [1, 2] and underwater exploration [3], there is increasing attention on processing and understanding underwater images. However, due to the unbalanced attenuation of underwater light [4], underwater images suffer from severe visual degradation, which often exhibits color casts, low contrast, and poor visibility [5, 6, 7]. There is an urgent need for UIE techniques to improve the visual quality of underwater images.
@@
Reviewing typical UIE strategies, early approaches based on physical models concentrate on accurately estimating medium transmission. This parameter represents the percentage of scene radiance reaching the camera [8, 9, 10]. Though physical model-based methods can achieve promising performance in some cases, they tend to produce unnatural and poor results in challenging underwater scenarios [11]. Recent approaches based on deep learning technology have gradually outperformed traditional physical model-based approaches [12, 13, 14]. Although simple regression-based convolutional networks may effectively enhance images in simple underwater scenarios, they often lack generalization and robustness when dealing with complex scenes. This limitation arises from their inability to fully model the degradation specific to real-world underwater conditions. Recently, image-to-image diffusion models have demonstrated remarkable success in effectively capturing intricate empirical distributions of images. These models facilitate the seamless transformation of images across diverse domains through iterative refinement steps, thereby presenting a new approach to address such challenging problems.
@@
However, image-to-image diffusion models heavily rely on paired datasets, and the inability to capture clear real-world reference images poses a challenge for solving the UIE problem [15, 16, 17]. To tackle this issue, researchers employed various approaches to synthesize multiple enhanced images from a single image and subsequently manually select the visually optimal one as a reference [4, 12]. Nevertheless, the synthesized paired datasets contain subjective preferences, which are manifested in the enhanced results, leading to the natural domain shift and weakened generalization. Regarding the unsupervised or semi-supervised strategies, they have also demonstrated the capability to generate visually captivating underwater images [18, 19], without necessitating paired datasets or relying solely on limited paired samples; however, their performance falls short when compared to that of the fully supervised method. Another alternative approach is to combine a pre-trained diffusion model, which captures the prior knowledge of mapping transitions between different image domains, with UIE benchmark datasets, thus reducing dependency on paired datasets [20]. However, fine-tuning or directly continuing training with the additional UIE benchmark datasets may potentially compromise valuable prior knowledge [21]. Furthermore, the lack of real image guidance in the reference domain limits the quality of enhancement results, leading to unnatural and unrealistic enhanced images.
@@
To address the above limitations, we propose CLIP-UIE, a new approach to treat UIE as a conditional underwater image generation problem. This approach is inspired by recent works on Denoising Diffusion Probabilistic Models [22, 23, 24, 25] and Contrastive Language-Image Pretraining (CLIP) guidance for image manipulation [26, 27, 28]. The proposed method is trained with a “denoising” objective to iteratively refine the enhancement result of the given input degraded underwater image. Initially, the key is to obtain a large number of underwater paired images. We utilize the image synthesis strategy based on color transfer [29, 30] to degrade in-air natural images into the real underwater domain, synthesizing the corresponding degraded underwater images, which is free of any subjective preferences. Subsequently, we train the basic image-to-image diffusion model on this large-scale synthetic paired dataset to learn the prior knowledge of mapping transitions from the underwater degradation domain to real in-air natural domain. This prior knowledge can counteract the unrealistic and unnatural synthetics effectively. It should be noted that the fine-tuning of the model to cater to downstream tasks is both inevitable and essential; however, this process may inadvertently erase the prior knowledge, resulting in overfitting and mode collapse [21]. Song et al. [31] and Jonathan et al. [32] have demonstrated that a pre-trained diffusion model can be effectively conditioned by leveraging the gradients of the classifier. Thus, we can exploit a trained classifier with the prior knowledge of in-air natural domain, combined it with UIE benchmark datasets to jointly fine-tune the diffusion model to avoid overfitting and mode collapse, bridging the domain gap between synthetic and real images.
Recently, it has been demonstrated that Vision-Language models encapsulate generic information, that can be paired with generative model to provide a simple and intuitive text-driven interface for image generation and manipulation [33], particularly the CLIP. The utilization of CLIP as a classifier to distinguish between in-air natural and underwater degraded images is somewhat effective; however, employing its gradients directly for controlling the reverse diffusion process remains challenging. Furthermore, locating precise prompts to accurately depict various forms of underwater degradation poses a significant challenge. Additionally, as discussed in [27], rephrasing similar prompts often results in substantial disparities in CLIP score. Similar issues for UIE problem are illustrated in Fig. 1. We train a CLIP-Classifier that contains the prior knowledge of the in-air natural domain by constraining text-image similarity in the CLIP embedding space. Moreover, for image enhancement tasks, we find that the image-to-image diffusion model and the CLIP-Classifier mainly act in the high-frequency region during the fine-tuning process. Therefore, we propose a fast fine-tuning strategy that focuses on the high-frequency range and can be up to 10 times faster. The contributions of this paper are summarized as follows:
@@
  • We propose employing color transfer techniques to simulate the degradation of $500$k in-air natural images of INaturalist [34] into underwater images, guided by an authentic underwater domain, thereby addressing the limited availability of extensive underwater image samples for enhancement tasks. The image-to-image diffusion model is then trained from scratch on this large-scale synthetic dataset to obtain pre-trained model.
@@
  • We propose a CLIP-Classifier that inherits the prior knowledge of the in-air natural domain, and then combine it with UIE benchmark datasets to jointly fine-tune the diffusion model to mitigate catastrophic forgetting and mode collapse. Experiments and ablation studies show that the proposed CLIP-UIE performs well and breaks the limitations of the reference domain to a certain extent.
@@
  • We find that for image enhancement tasks, which require consistent content, the image-to-image diffusion model and the CLIP-Classifier mainly act in high-frequency regions. Therefore, we propose a new fine-tuning strategy that only focuses on high-frequency regions, significantly improving the speed of fine-tuning, even up to 10 times.
@@
Figure 1: CLIP with proper prompts can serve as the classifier to evaluate the image quality and distinguish between in-air natural and underwater images. However, as a prevailing phenomenon, a simple approximate rephrasing of the prompt results in a significant change in CLIP score. On the contrary, our CLIP-Classifier with learned prompt (the last row of each example) shows more robust results for different types of in-air natural/underwater scenes.
@@ II-A Underwater Image Enhancement
In recent years, numerous approaches have been proposed in the field of underwater image enhancement, which can be classified into two distinct categories: traditional methods and deep learning-based techniques. For the traditional methods, researchers focus on directly adjusting image pixel values to produce a subjectively and visually appealing image, using techniques such as image fusion [35], histogram equalization [36], and pixel distribution [37], etc. Ancuti et al. [38] propose the fusion-based underwater image enhancement method, exploiting two images derived from a color-compensated and white-balanced version to prompt color contrast. However, these traditional methods are not robust to different types of underwater degradation [2], due to the omission of the underwater imaging mechanism. To improve the visual quality, physical model-based methods directly estimate the parameters of underwater imaging models based on prior information. For example, Fu et al. [8] utilize the retinex-based approach to capture a clean underwater image by combining enhanced reflectance and illumination. To further promote the performance, Zhuang et al. [10] improve the retinex-based algorithm by introducing Bayesian to estimate underwater medium transmissions. Nevertheless, owing to the scarcity of prior information, the enhanced image easily becomes unnatural and over/under-enhanced.
@@
Recently, deep learning has made significant strides in the fields of computer vision and image processing, leading to a series of endeavors aimed at enhancing the performance of underwater image enhancement through data-driven strategies. However, underwater images suffer from the unavailability of real reference images. Concretely, Li et al. [4] construct an underwater image benchmark (UIEB) by selecting the best result of various approaches as the reference. However, the performance of existing UIE methods, especially deep learning-based data-driven approaches, is limited by the unavoidable bias of those synthetic alternatives. Several GAN-based methods [18, 19] have been developed to generate pleasing underwater images, relaxing the requirement of paired underwater images. However, the performance of these methods lacks stability and generally falls behind fully supervised approaches, often introducing unrealistic elements and visual distortions.
@@
Utilizing in-air natural images to synthesize paired underwater degradation datasets provides a promising solution. Du et al. [20] transform in-air natural images into the pseudo underwater image domain by Koschmieder’s light scattering model, utilizing the prior knowledge of in-air natural images to counteract the adverse impact of unreal synthetic. However, the pseudo underwater domain based on the light scattering model sometimes cannot model the complexity of real-world underwater degradation. As a more intuitive and efficient alternative, we propose a new synthesis strategy based on color transfer [29] to generate paired datasets by degrading natural in-air images to the real underwater domain. This approach introduces additional prior knowledge: the mapping transitions between the real underwater degradation domain and the real in-air natural domain.
@@ II-B Diffusion Models
In recent studies, diffusion models have shown impressive performance in image generation by inverting image noising process using reverse-time stochastic differential equations (SDEs) [39]. However, this inversion process tends to alter the image context. To condition the denoising process, Classifier Guidance [32, 31] is an effective method to improve the sample quality of conditional diffusion models by utilizing the gradients of the classifier. An alternative strategy is to establish the relationship between input and conditional information in the latent space of the model [40, 25]. Controlling diffusion models facilitates personalization, customization, or task-specific image generation. Hence, the diffusion models have been successfully applied to various image enhancement tasks. The image-to-image diffusion model was first introduced by Saharia et al. [25] and has been exploited for image enhancement [41, 20], inpainting [42], and super-resolution [25], etc.
Conditional diffusion models have the ability to acquire mapping transitions between diverse image domains [21]. Building upon this observation, we train the conditional diffusion model on the synthetic dataset produced by color transfer to capture the prior knowledge of mapping transitions between the real underwater degradation domain and the real in-air natural domain. And then we will use a classifier to retain this valuable prior knowledge in subsequent fine-tuning of the pre-trained diffusion model, bridging the domain gap between synthetic and real images.
@@ II-C CLIP-based Image Manipulation
The CLIP model comprises a text encoder and an image encoder, which are jointly trained [43]. Leveraging the knowledge from 400 million carefully curated image-text pairs, it demonstrates the capability to accurately predict pairings of unseen (image, text) examples in a given batch. It has shown remarkable performance in text-to-image generation. Gal et al. [33] and Patashnik et al. [44] use CLIP models to modify StyleGAN-generated images with text prompts. More recently, Kim et al. [26] and Avrahami et al. [40] propose methodologies for leveraging the diffusion model guided by CLIP models in order to achieve global text-to-image synthesis, as well as facilitate local image editing. Text-driven generative models demonstrate strong performance when provided with explicit word prompts, such as “apple”, “banana”, and “orange”, which pertain to objective entities. However, it remains challenging to identify precise prompts for abstract image styles and visual qualities. Moreover, prompt engineering for abstract concepts necessitates the involvement of domain experts to meticulously annotate each image, and a notable challenge lies in the fact that similar prompts may yield disparate CLIP scores [28] (see Fig. 1). It restricts the user to generating and manipulating custom domains for abstract semantic concepts using the pre-trained CLIP.
To address the aforementioned issues, recent emerging research suggests fine-tuning the CLIP model with ample training data for domain adaptation [33]; however, this approach is labor-intensive and time-consuming. Concurrently, prompt learning with frozen CLIP model is the alternation to extract accurate low-level image representations [28], such as luminance, exposure, and contrast [27]. The present study also employs prompt learning to train a CLIP-Classifier, enabling the discrimination between in-air natural and underwater degraded images. Subsequently, the gradients derived from this CLIP-Classifier are employed to guide the sampling process of the diffusion model.
III Preliminary Knowledge and Proposed Method
III-A Preliminary Knowledge of Diffusion Models
Diffusion probabilistic models are a type of latent variable models that define a diffusion Markov chain to slowly introduce random noise into the data, and a corresponding reverse process to gradually remove the noise [22]. The diffusion process is a forward process that gradually injects noise into the data sample ${x}_{0}$ over $T$ time steps, diffusing the data samples through Gaussian transition: